Problem 15-2 Ayayai Company had the following stockholders equity as of January 1, 2017 Common stock, $5 par value, 18,200 shares issued Paid-in capital in excess of par-common stock Retained earnings $91,000 299,000 320,000 $710,000 Total stockholders equity During 2017, the following transactions occurred. Feb.1 Ayayai repurchased 1,990 shares of treasury stock at a price of $17 per share. Mar. 1 850 shares of treasury stock repurchased above were reissued at $15 per share. Mar. 18 530 shares of treasury stock repurchased above were reissued at $15 per share. Apr. 22 580 shares of treasury stock repurchased above were reissued at $19 per share. Your answer is partially correct. Try again. Prepare the journal entries to record the treasury stock transactions in 2017, assuming Ayayai uses the cost method. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation Debit Credit Feb. 1 ▼| Treasury Stock 33,830 Cash 33,830 Mar. 1 Cash 12,750 Retained Earnings 1,700 Treasury Stock 14,450 Mar. 18 Cash 7,950 Retained EarningsTreasury Stock Apr. 22 ▼ Cash 11,020 Treasury Stock Paid-in Capital from Treasury Stock SHOW LIST OF ACCOUNTS Your answer is partially correct. Try again. Prepare the stockholders equity section as of April 30, 2017. Net income for the first 4 months of 2017 was $126,200. (Enter account name only and do not provide descriptive information.) YAYAI COMPANY April 30, 2017 Common Stock 91,000 Paid-in Capital in Excess of Par - Common Stock 299,000 Paid-in Capital from Treasury Stock Toea Total Paid-in Capital Retained EarningsLess Treasury Stock Total Stockholders Equity Click if you would like to Show Work for this question: Open Show Work SHOW LIST OF ACCOUNTS
Answer #1

Where ever u have not filled i am filling them
Mar 18:
Cash(db) 530*15=7950
Retained earnings(db)1060
Treasury stock(cr) 530*17=9010
Apr 22:
Cash(db) 580*19=11020
Treasury stock(cr) 580*17=9860
Paid in capital from treasury stock(cr) 1160

stockholder equity section:

commn stock 91000
Paid in capital excess of par common stock 299000
Paid in capital from tresury stock 1160
Total paid in capital 391160
Retained earnings =320000-1060-1700+126200=443400
=391160+443400=834600
Less:
Treasury stock =33830-14450-9010-9860=510
Total stockholder equity 834090